OpenStudy (anonymous):

Oh shoot, I mistyped my previous problem :-/. f(x)=x^-1/2. Form and simplify into {f(x+h) - f(x)}/h

OpenStudy (anonymous):

\[f(x)= 1/\sqrt{x}\] just in case there's any confusion

OpenStudy (dumbcow):

this is definition of derivative using limits \[\large \lim_{h \rightarrow 0}\frac{\frac{1}{\sqrt{x+h}}-\frac{1}{\sqrt{x}}}{h}\] combine fractions on top, simplify into single fraction \[\large \lim_{h \rightarrow 0}\frac{\sqrt{x}-\sqrt{x+h}}{h \sqrt{x+h}\sqrt{x}}\] multiply top/bottom by conjugate \[\large \lim_{h \rightarrow 0}\frac{\sqrt{x}-\sqrt{x+h}}{h \sqrt{x+h}\sqrt{x}}*\frac{\sqrt{x}+\sqrt{x+h}}{\sqrt{x}+\sqrt{x+h}}\] \[\large \lim_{h \rightarrow 0}\frac{-h}{h \sqrt{x+h}\sqrt{x}(\sqrt{x}+\sqrt{x+h})}\] cancel the h's evaluate at h=0 \[= -\frac{1}{2x \sqrt{x}}\]